# LLM-FE: Automated Feature Engineering for Tabular Data with LLMs as Evolutionary Optimizers

### v1

@@ 3.3 Data-Driven Evaluation
Following hypotheses generation, we use the generated hypotheses to augment the original dataset with the newly derived features (Figure 1(b)). As illustrated in Figure 1(c), the feature space evaluation process comprises two stages: (i) model training on the augmented dataset, and (ii) performance assessment using validation data. We fit a tabular predictive model $f^{*}$, to the transformed training set $\mathcal{T}(\mathcal{X}_{\text{tr}})$, by minimizing the loss $\mathcal{L}_{f}$ as shown in Eq.2. Subsequently, we evaluated the LLM-generated feature transformations $\mathcal{T}$ by evaluating the model’s performance on the augmented validation set $\mathcal{T}(\mathcal{X}_{\text{val}})$ (see Eqs. 2 and 3). As explained in Section 3.1, the objective is to find optimal features that maximize the performance $\mathcal{E}$, i.e. accuracy for classification tasks and error metrics for regression problems.
3.4 Memory Management
@@
To promote diverse feature discovery and avoid stagnation in local optima, LLM-FE employs evolutionary multi-population memory management (Figure 1(d)) to store feature discovery programs in a dedicated database. Then, it uses samples from this database to construct in-context examples for LLM, facilitating the generation of novel hypotheses. This step consists of two components: (i) multi-population memory to maintain a long-term memory buffer, and (ii) sampling from this memory buffer to construct in-context example demonstrations. After evaluating the feature hypotheses in iteration $t$, we store the pair of hypotheses score ($\mathcal{T},s$) in the population buffer $\mathcal{P}_{t}$ to iteratively refine the search process. To effectively evolve a population of programs, we adopt a multi-population model inspired by the ‘island’ model employed by (Cranmer, 2023; Shojaee et al., 2024; Romera-Paredes et al., 2024). The program population is divided into $m$ independent islands, each evolving separately and initialized with a copy of the user’s initial example (see Figure 5(d)). This enables parallel exploration of the feature space, mitigating the risk of suboptimal solutions. At each iteration $t$, we select one of the $m$ islands and sample programs from the memory buffer to update the prompt with new in-context examples. The newly generated hypotheses samples $b$ are evaluated, and if their scores $s_{j}$ exceed the current best score, the hypotheses score pair ($\mathcal{T}_{j},s_{j}$) is added to the same island from which the in-context examples were sampled. To preserve diversity and ensure that programs with different performance characteristics are maintained in the buffer, we cluster programs within islands based on their signature, which is defined by their scores.
To build refinement prompts, we follow the sampling process from (Romera-Paredes et al., 2024), first sampling one of the $m$ available islands followed by sampling the $k$ programs from the selected island to create $k$-shot in-context examples for the LLM. Cluster selection prefers high-scoring programs and follows Boltzmann sampling (De La Maza & Tidor, 1992) with a score-based probability of choosing a cluster $i$: $P_{i}=\frac{exp(s_{i}/\tau_{c})}{\sum_{i}exp(s_{i}/\tau_{c})}$, where $s_{i}$ denotes the mean score of the $i$-th cluster and $\tau_{c}$ is the temperature parameter. The sampled feature transformation programs from the memory buffer are then included in the prompt as examples to guide LLM toward successful feature hypotheses. Refer to Appendix B.2 for more details.
@@
Algorithm 1 LLM-FE
0:  LLM $\pi_{\theta}$, Dataset $\mathcal{D}$, Metadata $\mathcal{M}$, Memory Buffer $\mathcal{P}$, Iterations $T$, Prediction model $f$, Performance Metric $\mathcal{E}$, Prompt template $p$
1:  # Initialize Memory Buffer
2:  $\mathcal{P}_{0}\leftarrow$ BufferInit()
3:  $-\mathcal{T}^{*},s^{*}\leftarrow\text{null},\text{-}\infty$
4:  # Update Prompt template with Data and Metadata $p\leftarrow\texttt{UpdatePrompt}(\mathcal{D,M})$
5:  for $t=1,\dots,T-1$ do
6:    # Add $k$ examples from Memory Buffer to the Prompt
7:    $p_{t}=(p+\mathcal{P}_{t-1}.\texttt{topk}())$
8:    # Sample $b$ outputs from the LLM
9:    $\mathcal{T}_{1},\dots,\mathcal{T}_{b}\leftarrow\pi_{\theta}(p_{t})$
10:    # Evaluation and Buffer Update
11:    for $j=0,\dots,b$ do
12:       $s_{j}\leftarrow$ FeatureScore($f,\mathcal{T}_{j},\mathcal{D},\mathcal{E}$)
13:       if $s>s^{*}$ then
14:          $\mathcal{T}^{*},s^{*}\leftarrow\mathcal{T}_{j},s_{j}$
15:          $\mathcal{P}_{t}\leftarrow\texttt{UpdateBuffer}(\mathcal{P}_{t-1},\mathcal{T}_{j},s_{j})$
16:       end if
17:    end for
18:  end for
18:  $\mathcal{T}^{*},s^{*}$
@@ 4 Experimental Setup
Table 2: Performance of XGBoost on Classification Datasets using various Feature Engineering (FE) Methods, evaluated using accuracy (higher values indicate better performance). We report the mean values and standard deviation across five splits. : denotes failure for all runs due to execution errors. bold: indicates the best performance. underline: indicates the second-best performance. ‘n’: indicates the number of samples; ‘p’: indicates the number of features.
@@
AutoFeat OpenFE CAAFE FeatLLM
balance-scale 625 4 0.856 $\pm$ 0.020 0.925 $\pm$ 0.036 0.986 $\pm$ 0.009 0.966 $\pm$ 0.029 0.800 $\pm$ 0.037 0.990 $\pm$ 0.013
breast-w 699 9 0.956 $\pm$ 0.012 0.956 $\pm$ 0.019 0.956 $\pm$ 0.014 0.960 $\pm$ 0.009 0.967 $\pm$ 0.015 0.970 $\pm$ 0.009
blood-transfusion 748 4 0.742 $\pm$ 0.012 0.738 $\pm$ 0.014 0.747 $\pm$ 0.025 0.749 $\pm$ 0.017 0.771 $\pm$ 0.016 0.751 $\pm$ 0.036
car 1728 6 0.995 $\pm$ 0.003 0.998 $\pm$ 0.003 0.998 $\pm$ 0.003 0.999 $\pm$ 0.001 0.808 $\pm$ 0.037 0.999 $\pm$ 0.001
cmc 1473 9 0.528 $\pm$ 0.029 0.505 $\pm$ 0.015 0.517 $\pm$ 0.007 0.524 $\pm$ 0.016 0.479 $\pm$ 0.015 0.531 $\pm$ 0.015
credit-g 1000 20 0.751 $\pm$ 0.019 0.757 $\pm$ 0.017 0.758 $\pm$ 0.017 0.751 $\pm$ 0.020 0.707 $\pm$ 0.034 0.766 $\pm$ 0.015
eucalyptus 736 19 0.655 $\pm$ 0.024 0.664 $\pm$ 0.028 0.663 $\pm$ 0.033 0.679 $\pm$ 0.024 0.668 $\pm$ 0.027
heart 918 11 0.858 $\pm$ 0.013 0.857 $\pm$ 0.021 0.854 $\pm$ 0.020 0.849 $\pm$ 0.023 0.865 $\pm$ 0.030 0.866 $\pm$ 0.021
pc1 1109 21 0.931 $\pm$ 0.004 0.931 $\pm$ 0.014 0.931 $\pm$ 0.009 0.929 $\pm$ 0.005 0.933 $\pm$ 0.007 0.935 $\pm$ 0.006
tic-tac-toe 958 9 0.998 $\pm$ 0.002 1.000 $\pm$ 0.000 0.994 $\pm$ 0.006 0.996 $\pm$ 0.003 0.653 $\pm$ 0.037 0.998 $\pm$ 0.005
vehicle 846 18 0.754 $\pm$ 0.016 0.788 $\pm$ 0.018 0.785 $\pm$ 0.008 0.771 $\pm$ 0.019 0.744 $\pm$ 0.035 0.761 $\pm$ 0.027
Mean Rank 3.18 3.09 3.00 3.82 1.54
Table 3: Performance of XGBoost on Regression Datasets using various Feature Engineering (FE) Methods, evaluated using normalized root mean square error (N-RMSE) (lower values indicate better performance). We report the mean values and standard deviation across five splits. : denotes failure for all runs due to execution errors. bold: indicates the best performance. underline: indicates the second-best performance. ‘n’: indicates the number of samples; ‘p’: indicates the number of features.
@@
Mean Rank 3.00 2.00 1.00
We evaluated LLM-FE on a range of tabular datasets, encompassing both classification and regression tasks. Our experimental analysis included quantitative comparisons with baselines and detailed ablation studies. Specifically, we assessed our approach using three known tabular predictive models with distinct architectures: (1) XGBoost, a tree-based model (Chen & Guestrin, 2016), (2) MLP, a neural model (Gorishniy et al., 2021), and (3) TabPFN (Hollmann et al., 2022), a transformer-based foundation model (Vaswani, 2017). The results highlight LLM-FE’s capability to generate effective features that consistently enhance the performance of different prediction models across diverse datasets.
4.1 Datasets
For our evaluation, we utilized two categories of datasets including (1) 11 classification datasets and (2) 10 regression datasets, each containing a mix of categorical and numerical features. Additionally, we included 8 large-scale, high-dimensional datasets to ensure a comprehensive evaluation of our method (see Appendix C.3). Following (Hollmann et al., 2024; Zhang et al., 2023), these datasets were sourced from well-known machine learning repositories, including OpenML (Vanschoren et al., 2014; Feurer et al., 2021), the UCI Machine Learning Repository (Asuncion et al., 2007), and Kaggle. Each dataset is accompanied by metadata, which includes a natural language description of the prediction task and descriptive feature names. We partitioned each dataset into training and testing sets using an 80-20 split. Following (Hollmann et al., 2024), we evaluated all methods over five iterations, each time using a distinct random seed and train-test splits. For more details, check Appendix 5.
4.2 Baselines
We evaluated LLM-FE against state-of-the-art feature engineering approaches, including OpenFE (Zhang et al., 2023) and AutoFeat (Horn et al., 2020), as well as LLM-based methods CAAFE (Hollmann et al., 2024) and FeatLLM (Han et al., 2024). We used XGBoost as the default tabular data prediction model in comparison between feature engineering baselines and employed GPT-3.5-Turbo as the default LLM backbone for all LLM-based methods. To ensure a fair comparison, all LLM-based baselines were configured to query the LLM backbone for a total of 20 samples until they converged to their best performance. Additional implementation details for all baselines are provided in Appendix B.1.
4.3 LLM-FE Configuration
@@ 5.1 Performance Comparisons
Evaluation on Classification Datasets.
In Table 2, we compare LLM-FE against various feature engineering baselines across 11 classification datasets. The results demonstrate that LLM-FE consistently enhances predictive performance from the base model (using raw data). LLM-FE also obtains the lowest mean rank, showing better effectiveness in enhancing feature discovery compared to other leading baselines. We have also extended our analysis to more complex datasets including high-dimensional and large-sample datasets to evaluate LLM-FE’s scalability in challenging scenarios. See Appendix C.3 for detailed results.
Evaluation on Regression Datasets.
To further evaluate the effectiveness of LLM-FE, we perform experiments on 10 regression datasets using the same evaluation settings employed for the classification datasets. The hypothesis space of current LLM-based baselines (CAAFE and FeatLLM) are only designed for classification datasets, so in this experiment, we only compare LLM-FE with non-LLM baselines (OpenFE and AutoFeat) which have been tested before for regression tasks. Table 3 compares the predictive performance using normalized root mean square error (N-RMSE). The results indicate that LLM-FE outperforms all baseline methods, achieving the lowest mean rank and consistently improving errors across all datasets.
@@ 5.2 Generalizability Analysis
To evaluate the generalizability of the LLM-FE framework, we examine its performance across multiple tabular prediction models and various LLM backbones. Specifically, we employ two LLM backbones, Llama-3.1-8B-Instruct and GPT-3.5-Turbo, in conjunction with three distinct tabular prediction models: XGBoost (Chen & Guestrin, 2016), a widely-used tree-based algorithm for tabular tasks; Multilayer Perceptron (MLP), a simple yet common deep-learning architecture tailored to tabular datasets (Gorishniy et al., 2021); and TabPFN (Hollmann et al., 2022), a recent transformer-based foundation model specifically designed for tabular data. Table 4 summarizes our findings, demonstrating that LLM-FE effectively identifies features that enhance the performance of various prediction models and LLM backbones across different tasks. Notably, the results indicate that features generated by LLM-FE using either LLM backbone consistently improve base model prediction performance compared to scenarios without any feature engineering. More detailed analyses of dataset-specific performances and feature transferability experiments are provided in Appendix C.1 and C.4.
Table 4: Performance improvement by LLM-FE using different prediction models and LLM backbones. We report the accuracy for classification tasks and normalized root mean square error for regression tasks. All results represent the mean and standard deviation computed across five splits. bold: indicates the best performance. TabPFN evaluations are conducted using only 10,000 samples due to its limited processing capacity.
@@
Base 0.820 $\pm$ 0.144 0.324 $\pm$ 0.016
LLM-FE Llama 3.1-8B 0.832 $\pm$ 0.147 0.310 $\pm$ 0.022
GPT-3.5 Turbo 0.840 $\pm$ 0.150 0.306 $\pm$ 0.015
@@
Base 0.745 $\pm$ 0.193 0.871 $\pm$ 0.027
LLM-FE Llama 3.1-8B 0.764 $\pm$ 0.173 0.794 $\pm$ 0.016
GPT-3.5 Turbo 0.784 $\pm$ 0.175 0.631 $\pm$ 0.043
@@
Base 0.852 $\pm$ 0.133 0.289 $\pm$ 0.016
LLM-FE Llama 3.1-8B 0.856 $\pm$ 0.133 0.288 $\pm$ 0.016
GPT-3.5 Turbo 0.863 $\pm$ 0.132 0.286 $\pm$ 0.015
Figure 2: Aggregated ablation study results across classification datasets, showcasing the impact of individual components on LLM-FE’s performance: (a) Data Examples, (b) Domain Knowledge, and (c) Evolutionary Refinement. The model performance is measured using accuracy, with higher accuracy indicating better performance.

### v2

@@ 3.3 Data-Driven Evaluation
As illustrated in Figure 1(b), we use the generated features to augment the original dataset with the newly derived features. Similar to hollmann2024large ; nam2024optimized , our feature evaluation process comprises two stages: (i) model training on the augmented dataset, and (ii) performance assessment for feature quality (Figure 1(c)). We fit a tabular predictive model $f^{*}$, to the transformed training set $\mathcal{T}(\mathcal{X}_{\text{tr}})$, by minimizing the loss $\mathcal{L}_{f}$ as shown in Eq.1. Subsequently, we evaluated the LLM-generated feature transformations $\mathcal{T}$ by evaluating the model’s performance on the augmented validation set $\mathcal{T}(\mathcal{X}_{\text{val}})$ (see Eqs. 1 and 2). As explained in Section 3.1, the objective is to find optimal features that maximize the performance $\mathcal{E}$, i.e., accuracy for classification and error metrics for regression.
3.4 Experience Management
Algorithm 1 LLM-FE
0:  LLM $\pi_{\theta}$, Dataset $\mathcal{D}$, Metadata $\mathcal{M}$, Iterations $T$, Model $f$, Metric $\mathcal{E}$
1:  $\mathcal{P}_{0}\leftarrow$ BufferInit()
2:  $\mathcal{T}^{*},s^{*}\leftarrow\text{null},-\infty$
3:  $p\leftarrow$ UpdatePrompt($\mathcal{D},\mathcal{M}$)
4:  for $t=1$ to $T{-}1$ do
5:     $p_{t}\leftarrow p+\mathcal{P}_{t-1}.\texttt{topk}()$
6:     $\{\mathcal{T}_{j}\}_{j=1}^{b}\leftarrow\pi_{\theta}(p_{t})$
7:     for $j=1$ to $b$ do
8:        $s_{j}\leftarrow$ FeatureScore($f,\mathcal{T}_{j},\mathcal{D},\mathcal{E}$)
9:        if $s_{j}>s^{*}$ then
10:           $\mathcal{T}^{*},s^{*}\leftarrow\mathcal{T}_{j},s_{j}$
11:        end if
12:        $\mathcal{P}_{t}\leftarrow$ UpdateBuffer($\mathcal{P}_{t-1},\mathcal{T}_{j},s_{j}$)
13:     end for
14:  end for
15:  return  $\mathcal{T}^{*},s^{*}$
@@ 3.4 Experience Management
To promote diverse feature discovery and avoid stagnation in local optima, LLM-FE employs evolutionary multi-population experience management (Figure 1(d)) to store feature discovery programs in a dedicated database. Then, it uses samples from this database to construct in-context examples for LLM, facilitating the generation of novel features. This step consists of two components: (i) multi-population memory to maintain a long-term memory buffer, and (ii) sampling from this memory buffer to construct in-context example demonstrations. After evaluating the feature transforms in iteration $t$, we store the pair of feature transforms and score ($\mathcal{T},s$) in the population buffer $\mathcal{P}_{t}$ to iteratively refine the search process. To effectively evolve a population of programs, we adopt a multi-population model inspired by the ‘island’ model employed by cranmer2023interpretable ; shojaee2024llm ; romera2024mathematical . The program population is divided into $m$ independent islands, each evolving separately and initialized with a copy of the user’s initial example (see Figure 6(d)). This enables parallel exploration of the feature space, mitigating the risk of suboptimal solutions. At each iteration $t$, we select one of the $m$ islands and sample programs from the memory buffer to update the prompt with new in-context examples. The newly generated feature samples $b$ are evaluated, and if their scores $s_{j}$ exceed the current best score, the feature score pair ($\mathcal{T}_{j},s_{j}$) is added to the same island from which the in-context examples were sampled. To preserve diversity and ensure that programs with different performance characteristics are maintained in the buffer, we cluster programs within islands based on their signature, defined by their scores. To build refinement prompts, we follow the sampling process from romera2024mathematical , first sampling one of the $m$ available islands, followed by sampling the $k$ programs from the selected island to create $k$-shot in-context examples for the LLM. Cluster selection prefers high-scoring programs and follows Boltzmann sampling de1992increased with a score-based probability of choosing a cluster $i$: $P_{i}=\frac{exp(s_{i}/\tau_{c})}{\sum_{i}exp(s_{i}/\tau_{c})}$, where $s_{i}$ denotes the mean score of the $i$-th cluster and $\tau_{c}$ is the temperature parameter. The sampled feature transformation programs from the memory buffer are then included in the prompt as examples to guide LLM toward successful feature transformations—incurring negligible computational overhead. Refer to Appendix 9 for more details.
@@ 4 Experimental Setup
We evaluated LLM-FE on a range of tabular datasets, encompassing classification and regression tasks. Our experimental analysis included quantitative comparisons with baselines and detailed ablation studies. Specifically, we assessed our approach using three known tabular predictive models with distinct architectures: (1) XGBoost, a tree-based model chen2016xgboost , (2) MLP, a neural model gorishniy2021revisiting , and (3) TabPFN hollmann2022tabpfn , a transformer-based foundation model vaswani2017attention . The results highlight LLM-FE’s capability to generate effective features that consistently enhance the performance of different prediction models across diverse datasets.
4.1 Datasets
We followed hollmann2024large to select datasets from han2024large ; hollmann2024large ; zhang2023openfe that include descriptive feature names and are not perfectly solvable using XGBoost. Our analysis contains 19 classification and 10 regression datasets, each containing mixed categorical and numerical features. We also include 8 large-scale, high-dimensional classification datasets to ensure comprehensive evaluation. These datasets were sourced from established machine learning repositories, including OpenML vanschoren2014openml ; feurer2021openml , UCI Machine Learning Repository asuncion2007uci , and Kaggle. Each dataset is accompanied by metadata, which includes a natural language description of the prediction task and descriptive feature names. We partitioned each dataset into train and test sets using an 80-20 split. Following hollmann2024large , we evaluated all methods over five iterations, each time using a distinct random seed and train-test splits. For more details, check Appendix 5.
4.2 Baselines
We evaluated LLM-FE against state-of-the-art feature engineering approaches, including OpenFE zhang2023openfe and AutoFeat horn2020autofeat , as well as LLM-based methods CAAFE hollmann2024large , FeatLLM han2024large and OCTree nam2024optimized . We used XGBoost as the default tabular data prediction model in comparison with baselines and employed GPT-3.5-Turbo as the default LLM backbone for all LLM-based methods. To ensure a fair comparison, all LLM-based baselines were configured to query the LLM backbone for a total of 20 samples until they converged to their best performance. Appendix B.1 contains additional implementation details.
4.3 LLM-FE Configuration
Table 2: Performance of XGBoost on Classification Datasets using various Feature Engineering (FE) Methods, evaluated using accuracy (higher values indicate better performance). We report the mean values and standard deviation across five splits. : denotes execution time of greater than 12 hours or failure due to execution errors. bold: indicates the best performance. underline: indicates the second-best performance. ‘n’: indicates the number of samples; ‘p’: indicates the number of features.
@@ 4.3 LLM-FE Configuration
AutoFeat OpenFE CAAFE FeatLLM OCTree
adult 48.8k 14 0.873 $\pm$ 0.002 0.873 $\pm$ 0.002 0.872 $\pm$ 0.002 0.842 $\pm$ 0.003 0.870 $\pm$ 0.002 0.874 $\pm$ 0.003
arrhythmia 452 279 0.657 $\pm$ 0.019 0.659 $\pm$ 0.018
balance-scale 625 4 0.856 $\pm$ 0.020 0.925 $\pm$ 0.036 0.986 $\pm$ 0.009 0.966 $\pm$ 0.029 0.800 $\pm$ 0.037 0.882 $\pm$ 0.022 0.990 $\pm$ 0.013
bank-marketing 45.2k 16 0.906 $\pm$ 0.003 0.908 $\pm$ 0.002 0.907 $\pm$ 0.002 0.907 $\pm$ 0.002 0.900 $\pm$ 0.002 0.907 $\pm$ 0.002
breast-w 699 9 0.956 $\pm$ 0.012 0.956 $\pm$ 0.019 0.956 $\pm$ 0.014 0.960 $\pm$ 0.009 0.967 $\pm$ 0.015 0.969 $\pm$ 0.009 0.970 $\pm$ 0.009
blood-transfusion 748 4 0.742 $\pm$ 0.012 0.738 $\pm$ 0.014 0.747 $\pm$ 0.025 0.749 $\pm$ 0.017 0.771 $\pm$ 0.016 0.755 $\pm$ 0.026 0.751 $\pm$ 0.036
car 1728 6 0.995 $\pm$ 0.003 0.998 $\pm$ 0.003 0.998 $\pm$ 0.003 0.999 $\pm$ 0.001 0.808 $\pm$ 0.037 0.995 $\pm$ 0.004 0.999 $\pm$ 0.001
cdc diabetes 253k 21 0.849 $\pm$ 0.001 0.849 $\pm$ 0.001 0.849 $\pm$ 0.001 0.849 $\pm$ 0.001 0.849 $\pm$ 0.001 0.849 $\pm$ 0.001
cmc 1473 9 0.528 $\pm$ 0.029 0.505 $\pm$ 0.015 0.517 $\pm$ 0.007 0.524 $\pm$ 0.016 0.479 $\pm$ 0.015 0.525 $\pm$ 0.027 0.531 $\pm$ 0.015
communities 1.9k 103 0.706 $\pm$ 0.016 0.704 $\pm$ 0.009 0.707 $\pm$ 0.013 0.593 $\pm$ 0.012 0.708 $\pm$ 0.016 0.711 $\pm$ 0.012
covtype 581k 54 0.870 $\pm$ 0.001 0.885 $\pm$ 0.007 0.872 $\pm$ 0.003 0.554 $\pm$ 0.001 0.832 $\pm$ 0.002 0.882 $\pm$ 0.003
credit-g 1000 20 0.751 $\pm$ 0.019 0.757 $\pm$ 0.017 0.758 $\pm$ 0.017 0.751 $\pm$ 0.020 0.707 $\pm$ 0.034 0.753 $\pm$ 0.021 0.766 $\pm$ 0.015
eucalyptus 736 19 0.655 $\pm$ 0.024 0.664 $\pm$ 0.028 0.663 $\pm$ 0.033 0.679 $\pm$ 0.024 0.658 $\pm$ 0.041 0.668 $\pm$ 0.027
heart 918 11 0.858 $\pm$ 0.013 0.857 $\pm$ 0.021 0.854 $\pm$ 0.020 0.849 $\pm$ 0.023 0.865 $\pm$ 0.030 0.852 $\pm$ 0.022 0.866 $\pm$ 0.021
jungle_chess 44.8k 6 0.869 $\pm$ 0.001 0.900 $\pm$ 0.004 0.901 $\pm$ 0.038 0.577 $\pm$ 0.002 0.869 $\pm$ 0.002 0.969 $\pm$ 0.004
myocardial 1.7k 111 0.784 $\pm$ 0.023 0.787 $\pm$ 0.026 0.789 $\pm$ 0.023 0.778 $\pm$ 0.023 0.787 $\pm$ 0.031 0.789 $\pm$ 0.023
pc1 1109 21 0.931 $\pm$ 0.004 0.931 $\pm$ 0.014 0.931 $\pm$ 0.009 0.929 $\pm$ 0.005 0.933 $\pm$ 0.007 0.934 $\pm$ 0.007 0.935 $\pm$ 0.006
tic-tac-toe 958 9 0.998 $\pm$ 0.002 1.000 $\pm$ 0.000 0.994 $\pm$ 0.006 0.996 $\pm$ 0.003 0.653 $\pm$ 0.037 0.997 $\pm$ 0.003 0.998 $\pm$ 0.005
vehicle 846 18 0.754 $\pm$ 0.016 0.788 $\pm$ 0.018 0.785 $\pm$ 0.008 0.771 $\pm$ 0.019 0.744 $\pm$ 0.035 0.753 $\pm$ 0.036 0.761 $\pm$ 0.027
Mean Rank 4.26 4.89 3.26 3.31 4.94 3.84 1.47
Table 3: Performance of XGBoost on Regression Datasets using various Feature Engineering (FE) Methods, evaluated using normalized root mean square error (N-RMSE) (lower values indicate better performance). We report the mean and standard deviation across five splits. bold: indicates the best performance. underline: indicates the second-best performance. ‘n’: indicates the number of samples; ‘p’: indicates the number of features.
@@
Mean Rank 3.40 3.10 2.20 1.00
@@ 5.1 Performance Comparisons
In Table 2, we compare LLM-FE against various feature engineering baselines across 19 classification datasets. The results demonstrate that LLM-FE consistently enhances predictive performance from the base model (using raw data). LLM-FE also obtains the lowest mean rank (best performance) at a lower computational cost (see Appendix 9), showing better effectiveness in enhancing feature discovery compared to other leading baselines. To further evaluate the effectiveness of LLM-FE, we perform experiments on 10 regression datasets using the same evaluation settings employed for the classification datasets. Due to the lack of regression data implementations in the available codebases for LLM-based baselines, in Table 3, we restrict our comparison to only non-LLM methods (OpenFE and AutoFeat), which have been previously validated on regression tasks. The results indicate that LLM-FE outperforms all baseline methods, achieving the lowest mean rank and consistently improving across all datasets.
@@ 5.2 Generalizability Analysis
Table 4: Performance improvement by LLM-FE using different prediction models and LLM backbones. We report the aggregated values for accuracy on classification tasks and normalized root mean square error on regression tasks. All results represent the mean and standard deviation computed across five splits. bold: indicates the best performance. TabPFN evaluations are conducted using only 10,000 samples due to its limited processing capacity.
@@
Base 0.820 $\pm$ 0.020 0.324 $\pm$ 0.016
LLM-FE Llama 3.1-8B 0.832 $\pm$ 0.021 0.310 $\pm$ 0.022
GPT-3.5 Turbo 0.840 $\pm$ 0.022 0.306 $\pm$ 0.015
@@
Base 0.745 $\pm$ 0.034 0.871 $\pm$ 0.027
LLM-FE Llama 3.1-8B 0.768 $\pm$ 0.032 0.794 $\pm$ 0.016
GPT-3.5 Turbo 0.791 $\pm$ 0.029 0.631 $\pm$ 0.043
@@
Base 0.852 $\pm$ 0.028 0.289 $\pm$ 0.016
LLM-FE Llama 3.1-8B 0.856 $\pm$ 0.017 0.288 $\pm$ 0.016
GPT-3.5 Turbo 0.863 $\pm$ 0.018 0.286 $\pm$ 0.015
To evaluate the generalizability of the LLM-FE, we examine its performance across multiple tabular prediction models and various LLM backbones. Specifically, we employ two LLM backbones, Llama-3.1-8B-Instruct and GPT-3.5-Turbo, in conjunction with three distinct tabular prediction models: XGBoost chen2016xgboost , a widely-used tree-based algorithm for tabular tasks; Multilayer Perceptron (MLP), a simple yet common deep-learning architecture tailored to tabular datasets gorishniy2021revisiting ; and TabPFN hollmann2022tabpfn , a recent transformer-based foundation model specifically designed for tabular data. Table 4 summarizes our findings, demonstrating that LLM-FE effectively identifies features that enhance the performance of various prediction models and LLM backbones across different tasks. Notably, the results indicate that features generated by LLM-FE using either LLM backbone consistently improve base model prediction performance compared to scenarios without any feature engineering. Detailed analyses on additional prediction models, hyperparameter optimization, feature transferability and more are provided in Appendix C.